# PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Spatiotemporal Prediction

### v2

@@ Abstract
Simulation of spatiotemporal systems governed by partial differential equations is widely applied in fields such as biology, chemistry, aerospace dynamics, and meteorology. Traditional numerical methods incur high computational costs due to the requirement of small time steps for accurate predictions. While machine learning has reduced these costs, long-term predictions remain challenged by error accumulation, particularly in scenarios with insufficient data or varying time scales, where stability and accuracy are compromised. Existing methods often neglect the effective utilization of multi-scale data, leading to suboptimal robustness in predictions. To address these issues, we propose a novel multi-scale learning framework, namely, the Physics-Informed Multi-Scale Recurrent Learning (PIMRL), to effectively leverage multi-scale data for spatiotemporal dynamics prediction. The PIMRL framework comprises two modules: the micro-scale module embeds physical knowledge into neural networks via pretraining, and the macro-scale module adopts a data-driven approach to learn the temporal evolution of physics in the latent space. Experimental results demonstrate that the PIMRL framework consistently achieves state-of-the-art performance across five benchmark datasets ranging from one to three dimensions, showing average improvements of over 9% in both RMSE and MAE evaluation metrics, with maximum enhancements reaching up to 80%.
1 Introduction
In the field of natural sciences, physical systems governed by partial differential equations (PDEs) have found widespread applications across disciplines including biology, chemistry, meteorology, etc. [1, 2, 3, 4, 5]. Although numerical methods have been regarded as reliable tools for modeling these systems, the use of Direct Numerical Simulation (DNS) faces significant hurdles due to inherent limitations. DNS necessitates high spatial resolution and fine time stepping, resulting in considerable computational demands and prolonged processing times. A case in point is the simulation of aerodynamic flows around aircraft, which typically requires the generation of millions of grid points, thereby imposing prohibitive computational requirements [6, 7]. Moreover, these numerical simulation methods require complete physical prior knowledge, such as PDE formula, parameters, and initial/boundary conditions [8].
The ongoing development of artificial intelligence (AI) has propelled research into data-driven simulation methods, showcasing significant potential [9, 10, 11, 12]. These methods, which do not require prior physics knowledge, offer user-friendly solutions. They also overcome traditional constraints on resolution and small time stepping, ensuring accurate solutions. Nevertheless, when faced with sparse data and multi-scale temporal challenges, these methods often struggle to optimally utilize data. They may either discard micro-scale data in favor of macro-scale data or vice versa, leading to compromised accuracy.
Multi-scale Burst sampling is a technique capturing multiple samples at a high rate over a short period of time to record rapidly changing events or transient phenomena, e.g., fast dynamics, followed by a low sampling rate to capture slow dynamics (see Figure 1). Compared to traditional continuous sampling methods, it provides high-resolution data at critical moments while maintaining resource efficiency. But conventional uniform-scale models struggle to fully utilize such data. The latent ODE approach [13] may work for multi-scale sampled data, but remains unclear whether it can address PDE problems. There is an urgent need for a new method to handle such a type of data for spatiotemporal systems.
@@ 1 Introduction
Data-driven methods face significant limitations when addressing challenges related to insufficient data and multiple time scales mentioned above. To overcome these obstacles, current approaches integrate physical knowledge into the model learning process. Specifically, physics-informed neural networks (PINNs) [14, 15, 16, 17, 18] design initial and boundary conditions as penalty terms in the loss function, thereby leveraging physical laws in a “soft” way. However, this soft embedding approach can sometimes lead to unsatisfactory results. Also, to some extent, the embedding methods used in PINNs are unable to ensure that the model fully adheres to the embedded physical conditions.
Directly embedding physical equations into the model architecture, as seen in the physics-encoded recurrent convolutional neural network (PeRCNN) [19], ensures strict adherence to the underlying physical laws. This approach addresses the issue of soft constraints in PINNs and enhances the model’s interpretability and generalization capabilities. However, this approach requires continuous micro-scale data, which imposes stringent demands on the dataset and can introduce instability in long-term predictions. Alternatively, methods such as the learned interpolation (LI) model [20] and multiscale simulation frameworks [21] integrate numerical techniques with neural networks. These hybrid approaches enhance simulation efficiency while maintaining acceptable accuracy, although they still have some of the limitations inherent in traditional numerical methods.
Moreover, we observe that the error increases along with the prediction horizon, a phenomenon that is widespread. Although existing efforts are predominantly placed on enhancing the model’s predictive capability, long-term prediction of spatiotemporal dynamics still suffers from error accumulation. Hence, properly controlling the accumulation of errors is becomes crucial. Based on this core idea, we are motivated to leverage comprehensive multi-scale data through a novel Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework to tackle the problems mentioned above. This framework consists of macro-scale and micro-scale modules. The micro-scale module, pretrained to learn the underlying physical laws, enhances the accuracy of simulations. The macro-scale module reduces the accumulation of errors by minimizing the number of rollout iterations for the micro-scale module, thereby enhancing long-term predictive performance. The main contributions of this paper are summarized as follows:
@@
  • We proposed a new PIMRL framework that effectively leverages information from multi-scale data for long-term spatiotemporal dynamics prediction. The concept of reducing the accumulation of errors is achieved through the integration of the macro-scale and micro-scale modules.
  • We designed a novel message passing mechanism between micro- and macro-scale modules, which effectively transmits physical information, enhances the micro-module’s correction capability, and reduces the number of micro-scale rollout iterations through the macro-scale module.
  • The PIMRL model achieved optimal performance in effectively predicting the duration of multiple different cases from fluid dynamics to physical systems, demonstrating its scalability and laying a solid foundation for more generalizable models in the future.
2 Related Work
Simulation tasks often aim to solve partial differential equations (PDEs) accurately and efficiently. Previous researchers have developed numerical methods achieving high precision with many nodes and short time steps. To speed up simulations, deep learning techniques have evolved from applying conventional algorithms to designing new models that integrate physical knowledge, including hybrid methods and models embedding physical principles.
Computational fluid dynamics. Computational Fluid Dynamics (CFD) is a branch of fluid mechanics that uses numerical methods and algorithms to analyze and solve problems involving fluid flow [1, 2, 3, 4, 5]. When faced with complex problems, unacceptable time and computational costs are the primary obstacles limiting the application of CFD. Moreover, when the corresponding physical equations contain unknown parameters or even unknown terms, numerical methods are unable to perform accurate simulations. The aforementioned methods have inspired us to leverage physical knowledge. However, to overcome the aforementioned issues, we have decided to introduce deep learning methods into our PIMRL framework.
Deep learning methods. With the advancement of AI, the application of AI in physical system simulation tasks has become more diverse and profound. For example, classical convolutional neural network models [11, 22] and ResNet [23], U-Net [24] what originally used for image segmentation, graph neural networks [25, 26], and transformer-based models [27, 28, 29, 30] are now being employed. Neural operators used for learning mappings between function spaces have also seen significant development in physical simulation tasks, such as DeepONet [9], MWT [31], FNO [10, 32, 33, 34], etc. In addition, there are several methods specifically designed for spatiotemporal prediction tasks, such as ConvLSTM [35], PredRNN [36], and TrajGRU [37]. The traditional deep learning methods mentioned above all require sufficient data for training; insufficient data can lead to poor model performance and severe overfitting issues. Moreover, error accumulation is another inevitable problem that limits the application of traditional deep learning methods. These limitations restrict the effectiveness and reliability of such models, especially in scenarios requiring long-term predictions or when data is scarce. The challenges outlined above have led us to propose a framework that effectively controls error accumulation and adeptly processes multi-scale data.
Physics-informed deep learning methods. To incorporate physical information into models, researchers have devised various methods. One category includes physics-inspired methods like PhyCRNet [38], PINN [14, 15, 16, 17, 18], PhyLSTM [39], PhySR [40], and PhyMPGN [41]. Another category involves physical embedding methods, where explicit physics knowledge is embedded into the model to fully leverage physical principles, like EquNN [42] and PDE-Net [43, 44]. PeRCNN [19, 45] offers a “hard” encoding mechanism to learn the dynamics of physical systems from limited data. This approach leverages prior physics knowledge for predictions, thereby equipping strong predictive capabilities and robust generalization across different initial conditions. However, all the aforementioned methods suffer from error accumulation, making it difficult to obtain stable and accurate results in long-term prediction tasks. PeRCNN effectively utilizes physical information in a way that can be leveraged within PIMRL.
Hybrid learning methods. In recent years, an emerging research direction has been to integrate deep learning methods with traditional numerical methods. By combining classical solvers with deep neural networks, hybrid approaches designed on this principle can achieve acceptable accuracy while operating faster than pure numerical methods. Examples include the Learned Interpolation (LI) model [20], numerical discretization learning [46], and P2C2Net [47]. However, in these hybrid methods, the numerical method component is not involved in the training process, and such approaches also require substantial amounts of data. The pioneering work multi-scale simulations of complex systems [21] employs a multi-scale framework for predictions, effectively reducing the required time and computational cost while maintaining good accuracy. However, this method still requires ample data for training. The idea behind hybrid learning methods is very promising, so we have introduced our own message-passing mechanism within our PIMRL framework.
3 Methodology
We propose the PIMRL framework, as illustrated in Figure 2, for spatiotemporal dynamics prediction with a small amount of multi-scale training data
3.1 Partial Differential Equations simulation task
Simulation tasks are intimately connected with PDEs, which are essential for describing and simulating physical models, and the time-dependent PDEs are defined as:
$\mathbf{u}_{t}=\mathcal{F}(t,x,\mathbf{u},\nabla\mathbf{u},\mathbf{u}\cdot\nabla\mathbf{u},\nabla^{2}\mathbf{u},\cdot\cdot\cdot;\mu),$ (1)
where $\mathbf{u}(x,t)$ denotes the spatiotemporal solution field, $\mathbf{u}_{t}$ the first-order time derivative, $\mathcal{F}(\cdot)$ a linear/nonlinear function, $\nabla$ the Nabla operator, $\nabla^{2}$ the Laplace operator, and $\mu$ the PDE parameter.
In addition, we define the initial and boundary conditions (ICs, BCs) for this equation, namely: $I[\mathbf{u}](x,t=0)=0$ and $B[\mathbf{u}](x,t)=0$ where $I$ and $B$ indicate IC and BC operators.
3.2 Overview
In this study, we aim establish a model capable of predicting the evolution of spatiotemporal dynamics for nonlinear PDE systems at a long-term horizon, given limited training data (e.g., a few trajectories resulted from different initial conditions). As outlined in the introduction, we have designed the PIMRL framework to mitigate error accumulation while ensuring that it can effectively capture the intrinsic changes in the physical system, rather than disregarding the physical information over large time intervals. Additionally, PIMRL is capable of efficiently utilizing multi-scale data. Below, we will provide a detailed exposition of the PIMRL framework’s architecture, the corresponding training methodologies, the macro and micro modules, as well as the boundary padding method that incorporates boundary conditions.
@@ 3.2 Overview
3.3 Forecasting Architecture
As shown in Figure 2, the PIMRL framework comprises two key components: the micro-scale module and the macro-scale module. The message-passing mechanism is an interaction between the micro-module and the macro-module. The information transmitted includes physical knowledge learned by the micro-module and corrections applied to the macro-module. Additionally, the macro-module also passes information to future iterations of the micro-module.
PIMRL is designed to achieve long-term prediction of spatiotemporal dynamics. PIMRL operates in a recursive manner, involving rollouts at the micro-module level, the macro-module level, and the overall framework level. Only the output from the macro-module contributes to the final output of the PIMRL framework and is used to compute the loss for training the entire framework, whereas the output from the micro-module serves as intermediate variables within the framework. And the message-passing mechanism is described as follows:
  • Firstly, the micro-module rollout is a simple autoregressive process with time step $\delta t$, where the output at the previous time step serves as the input for the next time step.
@@ 3.3 Forecasting Architecture
  • Finally, there is a PIMRL loop that operates in conjunction with the macro-module rollout. After every $N-1$ rollout of the macro-module, the micro-module stops participating in the prediction, and the macro-module performs $N$ steps of autoregressive prediction on its own. This completes a total of $2N$ rollouts. Each output from the macro-module during these $2N$ rollouts serves as the output of PIMRL as depicted in Equation 2.
We utilize two equations to represent the relationship between the micro-scale module and the macro-scale module, denoted as $F_{micro}$ and $F_{macro}$ respectively, as illustrated in Figure 2. The details of this process can be represented using the aforementioned symbols as follows:
[…]
@@
where the variable $\mathbf{u}$ denotes the physical state. The relationship between the micro-time step $\delta t$ and the macro-time step $\Delta t$ is given by $\Delta t=k\delta t$, where $k$ is an adjustable parameter determined by the time stepping of different scales in the real data.
3.4 Training Strategies
Since the PIMRL model consists of micro- and macro-scale modules, adopting brute-force end-to-end training yields unsatisfactory results (see the ablation study). Firstly, we establish a pre-training phase where only the micro-module is trained. The purpose of this pre-training phase is to enable the micro-module to effectively learn the dynamics of the physical system and the underlying physical laws, free from the influence of the macro-module. In the next phase, referred to as the overall training phase, all modules within the PIMRL framework are engaged in training together. During this phase, the micro-module benefits from the parameters pre-trained in the pre-training phase, serving to supervise and correct the macro-module. The output from the macro-module serves as the output of the entire PIMRL for computing the loss. The details are shown in Appendix.
3.5 Micro-scale Module
@@
The micro-scale module is designed to learn underlying physical laws that govern the spatiotemporal dynamics from micro-scale data with small time stepping, where we adopt the PeRCNN model [19] with the architecture of $\Pi$-block shown in Figure 3(a). In a forward Euler scheme: $\mathbf{u}_{(k+1)\delta t}=\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})\cdot\delta t+\mathbf{u}_{k\delta t},$ where $\delta t$ denotes that the module predicting in micro-scale time stepping. We can then approximate the $\mathcal{F}$ by $\hat{\mathcal{F}}$ described as follows:
@@
^$\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})=\sum^{N_{c}}_{c=1}W_{c}\cdot\left[\prod_{N_{l}}^{l=1}(K_{c,l}\star\hat{\mathbf{u}}_{k\delta t}+b_{l})\right].$ (5)
where $N_{c}$ denotes the channel count, and $N_{l}$ the total number of parallel convolutional layers. The symbol $\star$ denotes the convolutional operation. For each layer $l$ and channel $c$, $K_{c,l}$ designates the specific filter weight, while $b_{l}$ stands for the bias term of that layer $l$. In the context of a $1\times 1$ convolutional layer, $W_{c}$ denotes the weight assigned to the $c^{\text{th}}$ channel, with the bias term being omitted here for the sake of simplicity and brevity.
@@
When a certain term in the governing PDE remains known (e.g., the diffusion term $\Delta\mathbf{u}$), its discretization can be directly embedded in PeRCNN (called the physics-based Conv layer as shown in Figure 3(a)). The convolutional kernel in such a layer can be set according to the corresponding finite difference (FD) stencil. In essence, the physics-based Conv connection is constructed to incorporate known physical principles, whereas the $\Pi$-block is aimed at capturing the complementary unknown dynamics. The details of the physics-based FD Conv are provided in Appendix.
3.6 Macro-scale Module
The design of the macro-scale module, as a pivotal component of the PIMRL framework, is meticulously crafted to effectively manage and analyze macro-scale data. This type of data often poses unique challenges due to the substantial time spans it encompasses, which in turn lead to significant variations in the underlying physical states captured within the data points. These variations might be highly nonlinear and dynamic, making it difficult for traditional, physically motivated modeling methods to accurately capture all the nuances and complexities involved. As depicted in Figure 3, our macro-scale module utilizes ConvLSTM block with a residual connection. The structure of each block, illustrated in Figure 3(b), consists of a pair of encoders and decoders, along with a ConvLSTM cell shown in Appendix Figure. The input to this block undergoes mapping by the encoder, where the feature map serves as the characteristic of the latent space. Following this, the ConvLSTM cell simulates the dynamics, and the output is mapped back to the physical space through the decoder. Finally, the output, as a residual, is added to the feature from the previous time step to generate the prediction for the next time step. More details are provided in Appendix.
Figure 4: Periodic BC padding.
3.7 Boundary Condition Padding
Inspired by PeRCNN [19], we introduce BC hard encoding in both micro-scale and macro-scale modules. This encoding method ensures that the feature maps comply with the given BCs during the convolution process while also serving the purpose of padding, which involves filling the feature maps before convolution operations. Specifically, in this paper, our case adheres to periodic BCs, and the application of the corresponding padding is illustrated in Figure 4. This encoding scheme ingeniously incorporates BCs into the padding, thereby enhancing the accuracy of the prediction.
4 Experiments
To validate the effectiveness and versatility of our proposed PIMRL framework, we conducted extensive experiments on a diverse set of fluid dynamics and reaction-diffusion systems equations. Specifically, we tested our model on the following cases: the 1D Korteweg-de Vries (KdV) equation, the 2D Burgers equation, and three reaction-diffusion (RD) equations. These equations represent a range of physical phenomena with varying degrees of complexity and nonlinearity. Our results show that PIMRL consistently outperforms existing methods in terms of accuracy and robustness across these challenging cases.

### v4

# PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Burst-Sampled Spatiotemporal Dynamics

@@ Abstract
Deep learning has shown strong potential in modeling complex spatiotemporal dynamics. However, most existing methods depend on densely and uniformly sampled data, which is often unavailable in practice due to sensor and cost limitations. In many real-world settings, such as mobile sensing and physical experiments, data are burst-sampled with short high-frequency segments followed by long gaps, making it difficult to learn accurate dynamics from sparse observations. To address this issue, we propose Physics-Informed Multi-Scale Recurrent Learning (PIMRL), a novel framework specifically designed for burst-sampled spatiotemporal data. PIMRL combines macro-scale latent dynamics inference with micro-scale adaptive refinement guided by incomplete prior information from partial differential equations (PDEs). It further introduces a temporal message-passing mechanism to effectively propagate information across burst intervals. This multi-scale architecture enables PIMRL to model complex systems accurately even under severe data scarcity. We evaluate our approach on five benchmark datasets involving 1D to 3D multi-scale PDEs. The results show that PIMRL consistently outperforms state-of-the-art baselines, achieving substantial improvements and reducing errors by up to 80% in the most challenging settings, which demonstrates the clear advantage of our model. Our work demonstrates the effectiveness of physics-informed recurrent learning for accurate and efficient modeling of sparse spatiotemporal systems.
Introduction
@@
Learning dynamical systems from data is a central challenge in scientific machine learning. PDE-governed physical systems are ubiquitous in disciplines such as biology, chemistry, and meteorology (anderson1995computational; darwish2016finite). While traditional numerical solvers are reliable, direct numerical simulation (DNS) is often computationally prohibitive. It typically requires high spatial and temporal resolutions, long runtimes, and complete knowledge of the governing PDEs, model parameters, and initial/boundary conditions (ferziger2019computational; goc2021large).
To alleviate the high computational cost of DNS, data-driven deep learning methods such as DeepONet (lu2021learning), FNO (li2020fourier), and DPOT (hao2024dpot; azizzadenesheli2024neural) aim to learn solution operators directly from data. These approaches bypass the need for explicit PDEs and enable fast inference by learning mappings between function spaces. During rollout, they operate with relatively large time steps, thereby reducing simulation time significantly compared to classical solvers. However, their success critically depends on the availability of large, high-quality datasets, which are often expensive to obtain via numerical simulations or physical experiments (kanov2015johns; li2024learning).
To reduce data dependency, physics-informed learning has emerged as an effective paradigm. These methods embed physical knowledge into the training process either softly, by augmenting the loss function with PDE residuals, e.g., PINNs (raissi2019physics), PINO (li2024physics), MCNP (zhang2025monte), or more strictly, by designing model architectures that obey physical rules through discrete approximations, e.g., PeRCNN (Rao_2023; rao2022discovering), P2C2Net (WangEtAl2024NeurIPS), LDSolver (yan2025ldsolver), and TSM (sun2023neural). While these approaches improve physical fidelity and reduce labeled data requirements, they still face challenges in long-term forecasting. Due to their reliance on local update schemes, they typically require small time steps for numerical stability, which can lead to cumulative errors and increased computational cost over long horizons.
Since the performance of neural solvers is closely tied to the quality and quantity of training data, the strategy employed during data acquisition plays a critical role. However, most existing works assume that data are sampled uniformly or continuously over time. This assumption does not hold in many real-world scenarios, where data collection is often constrained by factors such as high experimental costs, limited temporal resolution, and hardware limitations. In practical applications including mobile sensing, in situ physical experiments, and computational fluid dynamics simulations, data are frequently gathered in a sparse and non-uniform manner, often across multiple temporal scales. Such irregular sampling patterns introduce information gaps and temporal inconsistencies, which can significantly degrade the performance of learning-based models that rely on temporally regular inputs (champion2019discovery).
Among alternative sampling strategies, burst sampling has recently emerged as a promising approach. It captures multiple high-frequency samples within short time intervals to record fast-changing or transient dynamics, while alternating with low-frequency or missing data periods to conserve resources. This scheme has demonstrated the ability to capture both rapid and slow dynamics efficiently (see Figure 1). However, despite its growing practical relevance, burst-sampled, multi-scale data remains largely unexplored in the context of deep learning for spatiotemporal systems.
To address this underexplored yet practically important challenge, we propose a framework named Physics-Informed Multi-Scale Recurrent Learning (PIMRL). PIMRL is designed to model sparse, non-uniform, and burst-sampled spatiotemporal data by coupling latent-space macro-scale inference with micro-scale physical correction. The macro module operates in a learned latent space to perform efficient temporal reasoning and accelerate simulation, while the micro module enforces fine-grained physical consistency using pretrained PDE priors. To integrate information across time and resolution scales, we introduce a dedicated cross-scale message passing mechanism that fuses information between the two modules, ensuring stability and accuracy during long-term rollouts. Our contributions are summarized as follows:
@@
  • We propose a Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework that effectively leverages information from multi-scale, burst-sampled data to improve long-term spatiotemporal dynamics prediction.
  • We develop a cross-scale message passing mechanism that fuses physical information between macro and micro modules, ensuring stable latent state evolution and mitigating error accumulation during long rollouts.
  • Our framework achieves optimal performance in effectively predicting the dynamics of diverse systems ranging from fluid flows to general physical phenomena, demonstrating scalability and generalizability under data-scarce, irregular sampling regimes.
Related Work
Accurate and efficient simulation of partial differential equations (PDEs) is essential for many scientific and engineering tasks, but traditional numerical methods are computationally costly. Although deep learning can accelerate simulations, these methods often require extensive physical knowledge or large datasets, resulting in poor performance when such resources are limited. In particular, current approaches struggle to model physical systems from sparse measurements without detailed physical priors.
Direct Numerical Simulation. Direct Numerical Simulation (DNS) is a computational method that solves the Navier–Stokes equations directly, capturing all spatial and temporal scales of turbulent flows without turbulence modeling (anderson1995computational; darwish2016finite). Although DNS yields highly accurate results, it is extremely computationally expensive. This immense computational cost and memory demand often limit DNS to relatively simple or small-scale problems, restricting its practical applicability. Furthermore, when the governing equations contain unknown parameters or missing physics, DNS may still struggle to provide reliable predictions. To overcome these challenges, PIMRL is specifically designed to efficiently predict outcomes by leveraging limited physical prior knowledge and data.
Deep Learning Methods. With the advancement of AI, its application to physical system simulation has become increasingly diverse and profound. For example, classical convolutional neural networks (CNNs) (bar2019learning), ResNet (lu2018beyond), graph neural networks (GNNs) (sanchez2020learning; pfaff2020learning), and transformer-based models (wu2024transolver; hang2024unisolver; janny2023eagle; li2024scalable) have been widely adopted. Neural operators for learning mappings between function spaces have also seen rapid development, such as DeepONet (lu2021learning), FNO (li2020fourier; rahman2022u; wen2022u), and PIANO (zhang2024deciphering).
Physics-Informed Deep Learning Methods. To incorporate physical knowledge, researchers have developed various physics-informed approaches, such as PhyCRNet (ren2022phycrnet), PINN (raissi2019physics), PeSANet (ijcai2025p862), and PDE-Net (long2018pde). These methods embed physical laws into deep models to improve prediction accuracy. PeRCNN (Rao_2023; rao2022discovering) employs a hard encoding of prior knowledge, enabling strong predictive and generalization capabilities even with limited data. However, most existing methods still suffer from error accumulation, which limits their long-term stability and accuracy. Notably, PeRCNN’s effective physical encoding strategy can be further leveraged in our proposed PIMRL framework.
Methodology
@@
Partial Differential Equations Simulation Task
Simulation tasks are closely related to PDEs, which fundamentally describe and model various physical phenomena, with the general form of a time-dependent PDE:
$\mathbf{u}_{t}=\mathcal{F}(t,x,\mathbf{u},\nabla\mathbf{u},\mathbf{u}\cdot\nabla\mathbf{u},\nabla^{2}\mathbf{u},\ldots;\mu),$ (1)
where $\mathbf{u}(x,t)$ denotes the spatiotemporal solution field, $\mathbf{u}_{t}$ is the first-order time derivative, $\mathcal{F}(\cdot)$ is a (possibly nonlinear) operator, $\nabla$ represents the gradient (nabla) operator, $\nabla^{2}$ is the Laplacian, and $\mu$ denotes the set of PDE parameters.
In addition, the initial and boundary conditions (ICs and BCs) are specified as, for example, $I[\mathbf{u}](x,t=0)=0$ and $B[\mathbf{u}](x,t)=0$, where $I[\cdot]$ and $B[\cdot]$ denote the operators for the initial and boundary conditions, respectively.
Overview
We aim to develop a model capable of predicting the evolution of spatiotemporal dynamics for nonlinear PDE systems over a long-term horizon, based on limited training data (e.g., a few trajectories generated from different initial conditions). As outlined in the introduction, we have designed the PIMRL to mitigate error accumulation and ensure that it effectively captures the intrinsic changes in the physical system, rather than overlooking significant physical information over extended time intervals. Furthermore, PIMRL is capable of efficiently utilizing multi-scale data.
As shown in Figure 2, PIMRL consists of two main components: a macro-scale module and a micro-scale module. The macro-module performs fast, coarse-grained inference in the latent space with large time steps, accelerating prediction and reducing error accumulation. The micro-module conducts fine-grained reasoning and adjusts the latent historical states. The message-passing mechanism enables interaction between these modules: the micro-module provides physical knowledge and correction signals to the macro-module, while the macro-module passes updated information to guide future micro-module iterations.
PIMRL operates in a recursive manner, involving rollouts at the micro-module level, the macro-module level, and the overall framework level. Only the output from the macro-module contributes to the final output of the PIMRL framework and is used to compute the loss for training the entire framework, whereas the output from the micro-module refines and updates the latent historical states. The message-passing mechanism we proposed is described as follows:
  • Firstly, the micro-module rollout is a simple autoregressive process with time step $\delta t$, where the output at the previous time step serves as the input for the next step.
@@
  • Finally, there is a PIMRL loop that operates in conjunction with the macro-module rollout. After every N1N-1 rollout of the macro-module, the micro-module stops participating in the prediction, and the macro-module performs $N$ steps of autoregressive prediction on its own. This completes a total of $2N$ rollouts. Each output from the macro-module during these $2N$ rollouts serves as the output of PIMRL as depicted in Equation 4.
Cross-scale Message Passing
We conceptualize the interaction between the micro-scale module ($F_{\text{micro}}$) and the macro-scale module ($F_{\text{macro}}$) as a cross-scale message passing framework, illustrated in Figure 2. The process is governed by the following operations:
First, during a data-rich burst, the micro-module generates a physically-informed “message” by performing a high-fidelity rollout over $k$ fine steps. This message generation process is defined by the Micro-module Loop:
$\text{Micro-module Loop: }\mathbf{u}^{\text{micro}}_{t+k\delta t}=\underbrace{F_{\text{micro}}(...F_{\text{micro}}}_{\times k}(\mathbf{u}_{t}))$ (2)
Next, this message, $\mathbf{u}^{\text{micro}}_{t+k\delta t}$, is passed to the macro-module, which consumes it to correct its own state. This state update is described by the Macro-module Loop:
$\text{Macro-module Loop: }\mathbf{\hat{u}}_{t+2k\delta t}=F_{\text{macro}}(\mathbf{u}^{\text{micro}}_{t+k\delta t})$ (3)
Finally, after being corrected by the message, the macro-module evolves autoregressively for efficient long-term forecasting. This complete, multi-step prediction cycle is summarized by the overall PIMRL process:
$\text{PIMRL: }\mathbf{\hat{u}}_{t+2N\Delta t}=\underbrace{F_{\text{macro}}(...F_{\text{macro}}}_{\times N}(\mathbf{\hat{u}}_{t+Nk\delta t}))$ (4)
@@
where the $\mathbf{u}$ denotes the physical state. The relationship between the micro-time step $\delta t$ and the macro-time step $\Delta t$ is given by $\Delta t=k\delta t$, where $k$ is an adjustable parameter.
Micro-scale Module
@@
The micro-scale module is designed to learn underlying physical laws that govern the spatiotemporal dynamics from micro-scale data with small time stepping, where we adopt the PeRCNN model (Rao_2023) with the architecture of $\Pi$-block shown in Figure 3(a). In a forward Euler scheme: $\mathbf{u}_{(k+1)\delta t}=\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})\cdot\delta t+\mathbf{u}_{k\delta t},$ where $\delta t$ denotes that the module predicting in micro-scale time stepping. We can then approximate the $\mathcal{F}$ by $\hat{\mathcal{F}}$ described as follows:
@@
$\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})=\sum^{N_{c}}_{c=1}W_{c}\cdot\left[\prod_{l=1}^{N_{l}}(K_{c,l}\star\hat{\mathbf{u}}_{k\delta t}+b_{l})\right],$ (5)
@@
where $N_{c}$ denotes the channel count, and $N_{l}$ the total number of parallel convolutional layers. The symbol $\star$ denotes the convolutional operation. For each layer $l$ and channel $c$, $K_{c,l}$ designates the specific filter weight, while $b_{l}$ stands for the bias term of that layer $l$. In the context of a $1\times 1$ convolutional layer, $W_{c}$ denotes the weight assigned to the $c^{\text{th}}$ channel, with the bias term being omitted for the sake of simplicity and brevity. When a certain term in the governing PDE remains known (e.g., the Laplace operator $\nabla^{2}\mathbf{u}$), its discretization can be directly embedded in PeRCNN (called the physics-based Conv layer as shown in Figure 3(a)). The convolutional kernel in such a layer can be set according to the corresponding finite difference (FD) stencil. In essence, the physics-based Conv connection is constructed to incorporate known physical principles, whereas the $\Pi$-block is aimed at capturing the complementary unknown dynamics. The details of the physics-based FD Conv are provided in the Physical Filter part of Appendix A.
Macro-scale Module
The macro-scale module serves as the long-term predictive module of PIMRL, specifically engineered to bridge the large temporal gaps inherent in burst-sampled data. Instead of operating in the computationally expensive physical space, it performs inference within a compact latent space. This design choice is motivated by the need for both computational efficiency and the expressive power to capture complex, long-range dependencies that are intractable for traditional, fine-grained solvers.
The module is implemented as a residual ConvLSTM autoencoder (Figure 3; Appendix B), with components operating together within a message-passing framework. An encoder maps the input to a latent space, where the ConvLSTM cell propagates the system’s state forward in time with large, efficient steps ($\Delta t$). Crucially, this latent-space evolution is not a pure black-box rollout. It is periodically corrected by physics-informed messages from the micro-module. This mechanism allows the macro-module to anchor its latent trajectory to physical reality, preventing the error accumulation and drift that plague purely data-driven recurrent models during long-term forecasting. By integrating the efficiency of latent-space dynamics with the physical fidelity of micro-scale corrections, our macro-scale module achieves a robust and accurate forecasting capability.
Case Numerical Methods Spatial Grid Macro Step Micro Step Training Trajectories Test Trajectories
KdV FVM 256 0.01s 0.01s 5 2
Burgers FDM $128^{2}$ 0.001s 0.01s 13 3
FN FDM $128^{2}$ 0.5s 0.01s 5 3
2D GS FDM $128^{2}$ 0.002s 0.01s 2 3
3D GS FDM $96^{3}$ 0.25s 0.01s 3 2
Table 1: Summary of experimental settings for different cases. The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training.
Experiments
To validate the effectiveness of our proposed PIMRL framework, we conducted extensive experiments on a diverse set of fluid dynamics and reaction-diffusion systems.
Figure 4: Error propagation curves and final prediction plots for the KdV case, comparing the PIMRL and baseline models.
Figure 5: Figures (a)–(c) summarize qualitative comparisons with baselines on Burgers, FN, and 2D GS, showing error propagation and final predictions.